# Understanding World or Predicting Future? A Comprehensive Survey of World Models

### v2

@@ 5.1.2. World Simulators
5.2. Robots
World models have emerged as a transformative paradigm in robotics, empowering robots to perceive, predict, and act effectively in complex environments. This progress is driven in part by advances in neural architectures (vaswani2017attention, ; ho2020denoising, ) and learning algorithms (schulman2017proximal, ; rafailov2024direct, ), which enable robots to build implicit representations that capture key aspects of the external world. Complementarily, prediction models (finn2016unsupervised, ; finn2017deep, ) offer the ability to forecast future environmental states, moving beyond static abstractions to support anticipatory and adaptive behavior. Together, these capabilities make it increasingly feasible for robots to learn directly from real-world interactions. In Table 5, we summarize the core learning tasks involved in constructing world models for robotics, categorized according to the three major perspectives outlined above (typical examples shown in Figure S1).
5.2.1. Learning Implicit Representation
Task Model Year Input Backbone
ViT (dosovitskiy2020image, ) 2020 Image Transformer
RoboCraft (shi2024robocraft, ) 2024 Image GNN
Predictive Coding (gornet2024automated, ) 2024 Image ResNet
SpatialLM (spatiallm, ) 2025 3D point clouds MLLM
Task Representation BC-Z (jang2022bc, ) 2022 Text& Video LLM& ResNet
Text2Motion (lin2023text2motion, ) 2023 Text LLM
Gensim (wang2023gensim, ) 2023 Text LLM
VIPER (escontrela2024video, ) 2024 Video Transformer
GR-2 (cheang2024gr, ) 2024 Text & Video Transformer
IRASim (zhu2024irasim, ) 2024 Trajectory Diffusion
SWIM (mendonca2023structured, ) 2023 Video Transfer Learning
CoSTAR (agha2021nebula, ) 2021 Multimodal Belief Space
Evaluation OpenEQA (majumdar2024openeqa, ) 2024 Image& Text LLM
5.2.2. Predicting Future States of the Environment
Robotic tasks are inherently sequential and long-term, where early decisions significantly influence future outcomes (simmons1994structured, ). Effective robotic planning thus relies heavily on accurately predicting future environmental states, enabling proactive decision-making and reducing costly errors. Classic closed-loop algorithms (bertsekas2012dynamic, ; khatib1987unified, ), which select actions solely based on current observations, are typically short-sighted, risking irreversible mistakes. Additionally, approaches depending on explicit dynamic models crafted from expert knowledge tend to be limited in flexibility and robustness.
A key recent insight is the use of generative video models—particularly those leveraging diffusion (esser2023structure, ; chi2023diffusion, ; black2023zero, ; helearning, ) and transformer architectures (yu2023magvit, ; yan2021videogpt, )—to implicitly learn environmental dynamics directly from visual data. For instance, UniPi (du2024learning, ) frames action prediction explicitly as a video generation problem, conditioning a constrained diffusion model on the current state to visualize future scenarios. Similarly, VIPER (escontrela2024video, ) employs a pretrained autoregressive transformer to guide robotic actions, effectively leveraging rich representations learned from expert demonstration videos. IRASim (zhu2024irasim, ) leverages diffusion models for trajectory-to-video generation tasks, starting from an initial given frame. Moreover, models like GR-2 (wu2023unleashing, ; cheang2024gr, ) benefit from the vast scale of internet videos to establish robust priors, subsequently fine-tuning on specific robotics tasks to generate accurate image predictions and action trajectories.
Collectively, these methods demonstrate the promise of generative, vision-centric modeling as a foundation for anticipatory robotic control, significantly enhancing robots’ capabilities to reason (to some extent) about future states and improve long-term task performance.
5.2.3. From Simulation to Real World
5.3. Social Simulacra
The concept of “social simulacra” was originally introduced as a prototyping technique in (park2022social, ), aimed at helping designers create a virtual social computing system encompassing many diverse agents. The traditional methods of constructing agents based on expert-defined rules (schelling1971dynamic, ; brock1998heterogeneous, ) or reinforcement learning (zheng2022ai, ) face issues such as overly simplistic behaviors or a lack of interpretability. However, the emergence of LLMs provides a transformative tool for building more realistic social simulacra, achieving more convincing stylized facts (li2024econagent, ) or accurate predictions. Social simulacra can be seen as a form of world model that mirrors real-world social computing systems. From another perspective, the agents within social simulacra also develop implicit representations of the external system; that is, they build an implicit world model that supports the generation of their social behaviors. The relationship between the world model and social simulacra is shown in Figure S2, and the summary of representative works is shown in Table 6.
Table 6. Representative works of social simulacra.
Advance What to simulate Effects of world model
@@ 5.3. Social Simulacra
World Model as Social Simulacra AI Town (park2023generative, ) Machine society Stylized facts
S3 (gao2023s, ) Social network Predictions
Papachristou et al. (papachristou2024network, ) Social network Stylized facts & Predictions
Xu et al. (xu2023exploring, ) Games Stylized facts
@@
Project Sid (al2024project, ) Collective rules (tax) Stylized facts
AgentSociety (piao2025agentsociety, ) Social life Stylized facts
@@
World Model in Social Siumlacra Agent-Pro (zhang2024agent, ) Games Belief
Zhang et al. (zhang2023exploring, ) Machine society Reflection & Debate
GovSim (piatti2024cooperate, ) Resource sharing Cognition
AgentGroupChat (gu2024agent, ) Group chat Belief & Memory
5.3.1. Building Social Simulacra Mirroring Real-world Society
In the era of the rapid rise of LLM agents, building realistic social simulation systems becomes more practical. One of the most famous examples of social simulacra is AI Town (park2023generative, ), a world model composed of 25 generative agents, essentially forming a sandbox social environment. In this virtual community, the agents exhibit believable individual behaviors, and at the group level, emergent social behaviors similar to those that might appear in the real world. Along these lines, more and more attempts are being made to replace humans in various social scenarios with LLM agents, in effect forming their own scenario-specific social simulacra. These works have used the simulacra paradigm in such scenarios as social networks and cooperative or competitive games, among others (gao2024large, ).
Recent studies have extended social simulacra by demonstrating how LLM agents can reflect realistic social interactions and collective behaviors, effectively mirroring human societies. A critical insight is that LLM-driven agents, by modeling aspects such as emotions, attitudes, and decision-making patterns, can reproduce intricate social phenomena across multiple contexts. For example, S3 (gao2023s, ) leverages these human-like features to simulate realistic message propagation on social networks, successfully capturing the dynamics of real-world public events. Related studies (papachristou2024network, ) deepen this understanding by analyzing how network structures emerge organically among LLM agents, comparing these structures directly to those formed in human social networks. Beyond networks, the capacity of LLM agents to embody sophisticated strategic reasoning is highlighted in simulations of social games such as Werewolf (xu2023exploring, ), where the agent outputs exhibit patterns resembling human strategic behaviors such as deception and confrontation. Furthermore, in economic contexts, works such as EconAgent (li2024econagent, ) reveal that collective behaviors emerging from individual economic reasoning by LLM agents can convincingly reproduce macroeconomic trends and regularities. Recently, AgentSociety (piao2025agentsociety, ) has further expanded this paradigm by creating a large-scale, LLM-powered societal simulation capable of modeling diverse social phenomena such as polarization and response to public policies, thus providing a versatile platform for computational social science research. These studies collectively illustrate the broad potential of social simulacra, demonstrating that agent-based world models powered by LLMs can capture diverse and realistic social, strategic, and economic interactions.
@@ 5.3.2. Agent’s Understanding of External World in Social Simulacra
LLM agents build their memory by storing observations obtained through interactions with the external environment (zhang2024survey, ), thereby forming implicit representations and basic cognition of the external world, especially in the context of simulating social scenarios. This cognition is stored in a memory bank in textual form for LLM agents to retrieve and use, enabling them to access useful information and fully leverage experiential knowledge from past interactions with the environment when making decisions.
Agent-Pro (zhang2024agent, ) transforms the memory of its interactions with the external environment (specifically with other agents in interactive tasks) into what are called ‘beliefs’. Based on these beliefs, it makes the next decision and updates its behavior strategy. These beliefs represent the agent’s social understanding of the environment and other agents within it, relating to Theory of Mind mentioned in Section 3.2. Other works on LLM agents have also adopted similar designs. For example, Zhang et al. (zhang2023exploring, ) introduces mechanisms of reflection and debate from a social psychology view for modeling multi-agent collaboration tasks. A more advanced study is GovSim (piatti2024cooperate, ), which explores whether cooperative behaviors aimed at sustainable resource development can emerge within a society composed of LLM agents. In this setup, each agent gathers information about the external world and other agents’ behavioral strategies through multi-agent conversations and subsequently forms its own high-level insights, essentially creating an implicit representation of the world model. Another similar application scenario is Interactive Group Chat (gu2024agent, ), where human-like behaviors and strategies emerge across four narrative scenarios, including Inheritance Disputes, Law Court Debates, etc.
@@ 6.1. Physical Rules and Counterfactual Simulation
A key objective of world models is to capture the causal structure of their environments—especially the underlying physical rules—so they can reason about counterfactuals beyond the data distribution (pearl2009causal, ). This capacity is crucial for handling rare, mission-critical events (e.g., autonomous-driving corner cases (feng2023dense, )) and for narrowing sim-to-real gaps. Recent progress raises the question of whether large-scale, purely data-driven generative models can acquire such rules from raw visual data alone. While transformer- and diffusion-based video generators such as Sora (sora2024, ) produce strikingly realistic sequences, studies reveal persistent physical-law failures—e.g., inaccurate gravity, fluid, or thermal dynamics (wang2023newton, ).
Hybrid approaches that explicitly embed physics are emerging as promising alternatives. Genesis(Genesis, ) illustrates this direction by unifying fast, photo-realistic rendering with a re-engineered universal physics core, allowing language-conditioned data generation grounded in first-principles simulation. PhysGen(liu2024physgen, ) takes a similar stance at the image-to-video level: it couples a rigid-body simulator with a diffusion refiner, enabling controllable, physically plausible motion from a single image. Complementary diagnostic work underscores why such hybrids are needed. Kang et al.(kang2024far, ) show that scaling diffusion video models yields perfect in-distribution fidelity yet breaks down on out-of-distribution or combinatorial tests, indicating “case-based” rather than rule-based generalization. Motamed et al.(motamed2025generative, ) reach a similar conclusion with the Physics-IQ benchmark: current video generators achieve visual realism but largely fail on tasks requiring understanding of optics, fluid dynamics, or magnetism.
Taken together, the evidence suggests that data-driven scaling alone is insufficient to recover robust physical laws. Integrating explicit simulators—or otherwise enforcing physical priors (shi2022learning, )—remains a promising path toward world models that generalize to unseen counterfactual scenarios while retaining interpretability and transparency.

### v4

@@ 5.3.1. Autonomous Driving
Table 6. Applications of World Models in Autonomous Logistics and Urban Analytics
Category Sub-category Paper Year Venue
Autonomous Logistics Micromobility NWM (bar2025navigation, ) 2025 CVPR
URBAN-SIM (wu2025towards, ) 2025 CVPR
Vid2Sim (xie2025vid2sim, ) 2025 CVPR
CityWalker (liu2025citywalker, ) 2025 CVPR
Aerial AirScape (zhao2025airscape, ) 2025 ACM MM
CityNavAgent (zhang2025citynavagent, ) 2025 ACL
CityEQA (zhao2025cityeqa, ) 2025 EMNLP
UrbanVideo-Bench (zhao2025urbanvideo, ) 2025 ACL
Urban Analytics Knowledge CityGPT (feng2024citygpt, ) 2025 KDD
UrbanLLaVA (feng2025urbanllava, ) 2025 ICCV
GeoLLM (manvi2023geollm, ) 2024 ICLR
Prediction GPS-to-Image (feng2025gps, ) 2025 CVPR
AI’s Blind Spots (beneduce2025ai, ) 2025 arXiv
Understanding AgentMove (feng2025agentmove, ) 2025 NAACL
CAMS (du2025cams, ) 2025 arXiv
PIGEON (lan2025open, ) 2025 ACL
5.3.2. Autonomous Logistics
This section introduces the application of world models in the autonomous logistics for urban scenario, primarily focusing on two areas: 1) miniature mobile logistics vehicles and 2) low-altitude aerial vehicles. In the research of both domains, we will introduce specific work from the perspectives of both understanding and prediction, which are shown in Table 6.
In the context of miniature mobile logistics vehicles, which represent an extension of autonomous driving into embodied intelligence scenarios, there is a need to cope with more complex surrounding environments and human-robot interactions. The core task is navigation, which involves comprehending the surrounding world to achieve more efficient and safer movement and interaction. In terms of understanding, studies like Vid2Sim (xie2025vid2sim, ) and CityWalker (liu2025citywalker, ) leverage the vast amount of videos available on the internet to understand diverse environments and interaction scenarios among movement. This understanding is then used to train the robot’s navigation policy, aiming for strong generalization and controllability. In terms of predicting the future, two typical paradigms exist. The first involves constructing a realistic physical simulator (dosovitskiy2017carla, ; wu2025towards, ) to generate diverse scenarios. Through large-scale training in this virtual environment, the robot’s capabilities are enhanced and then applied to the real world. The other paradigm is based on an interactive and controllable world model (bar2025navigation, ) derived from video generation models. This model generates potential future scenes that the robot might encounter based on its actions, specifically its trajectory, thereby enhancing its generalized navigation capabilities in a variety of scenarios.
Regarding low-altitude aerial vehicles, the vast majority of current work focuses on applications in understanding, particularly in scene comprehension and navigation (zhao2025cityeqa, ; zhao2025urbanvideo, ; zhang2025citynavagent, ), while applications in prediction (zhao2025airscape, ) are still in their nascent stages. In the realm of understanding, a typical paradigm involves comprehending the current urban scene and its key elements from images (zhao2025cityeqa, ) or videos (zhao2025urbanvideo, ). This provides sufficient support for subsequent tasks such as navigation (zhang2025citynavagent, ). Understanding diverse and complex urban scenes relies on strong prior knowledge of the urban environment, especially the rich world knowledge and common sense derived from large language models (achiam2023gpt4, ). In the domain of generation, AirScape (zhao2025airscape, ) stands as the first world model for low-altitude aerial vehicles. It enables the prediction of future scenes based on the aerial’s actions while maintaining physical and spatiotemporal consistency. This provides a new approach and environment for the efficient training and task resolution of low-altitude aerial vehicles in the future.
5.3.3. Urban Analytics
In this section, we primarily introduce the application of world models in Urban Analytics. Building on the premise mentioned in Section 3.2 that large language models have already learned worldly geographical knowledge (feng2024citygpt, ; manvi2023geollm, ), we will introduce relevant work from the perspectives of both understanding and prediction. Related works are summarized in Table 6.
In terms of understanding, on one hand, concerning the environment, multimodal large language models such as UrbanLLaVA (feng2025urbanllava, ) leverage the rich world knowledge embedded within models to perform generalizable tasks like urban scene recognition and comprehension (feng2024citybench, ). On the other hand, regarding human behavior within the environment, AgentMove (feng2025agentmove, ) and CAMS (du2025cams, ) utilize existing or specifically enhanced urban geospatial knowledge within models (feng2024citygpt, ) to model human mobility patterns. Meanwhile, PIGEON (lan2025benchmarking, ) draws upon the common sense knowledge of large language models to understand daily human needs and their corresponding behaviors, thereby achieving accurate understanding and prediction even for infrequent scenarios.
In terms of prediction, GPS-to-Image (feng2025gps, ) attempts to use GPS signals to control the style and scene of generated images, demonstrating the feasibility of using models to grasp the relationship between geographical location and urban scenery. However, researchers (beneduce2025ai, ) have further discovered that existing image generation models still have significant room for improvement in accurately distinguishing the cultural styles and scenic characteristics of different geographical locations.
Overall, the application of world models in urban analytics is still relatively limited, indicating substantial potential for future applications.
5.4. Societal Intelligence
Societal intelligence is the collective ability of a society to sense its environment, reason about possible futures, and coordinate actions toward shared goals (levy1997collective, ; malone2015handbook, ; malone2018superminds, ). It emerges from the interactions among individuals, institutions, and their surroundings (epstein2012generative, ). One effective way to operationalize societal intelligence in silico is through social simulacra (park2022social, ), which are virtual social computing systems populated by diverse agents capable of realistic, human-like behaviors. Traditionally, such systems have been constructed using expert-defined rules (schelling1971dynamic, ; brock1998heterogeneous, ), which encode domain knowledge into explicit behavior specifications, or through reinforcement learning (zheng2022ai, ), which optimizes agent strategies via trial-and-error in simulated environments. While effective in certain contexts, these approaches often lead to overly simplistic dynamics or limited interpretability. The advent of LLMs offers a transformative foundation for creating richer and more convincing simulacra, enabling both the reproduction of stylized facts (li2024econagent, ) and the generation of credible predictions (piao2025agentsociety, ).
In this survey, we examine world models within the framework of societal intelligence from two complementary perspectives. First, social simulacra can function as an explicit world model that mirrors real-world societies, providing a structured environment in which societal intelligence can emerge. Second, agents in simulacra can develop implicit world models through interaction, forming internal representations of the external environment that guide their decisions and social behaviors. These two perspectives—mirroring real-world society and understanding the external world—structure the following subsections. Table 7 summarizes representative works across both perspectives.
Table 7. Representative works of LLM-driven social simulacra from two perspectives: mirroring real-world society and understanding the external world.
Representative work Simulation focus Role of world model
@@ 5.4. Societal Intelligence
Mirroring Real-world Society Generative Agents (park2023generative, ) Daily social life Stylized facts
AI Town (park2023generative, ) Sandbox community Stylized facts
S3 (gao2023s, ) Social networks Predictions
Papachristou et al. (papachristou2024network, ) Network formation Stylized facts
Xu et al. (xu2023exploring, ) Social games Strategic patterns
@@
Project Sid (al2024project, ) Collective rules Stylized facts
OASIS (yangoasis, ) Social media Stylized facts
GenSim (tanggensim, ) Social media Stylized facts
YuLan-OneSim (wang2025yulan, ) General platform Stylized facts
AgentSociety (piao2025agentsociety, ) General platform Stylized facts & predictions
SocioVerse (zhang2025socioverse, ) General platform Stylized facts & predictions
@@
Understanding the External World Agent-Pro (zhang2024agent, ) Interactive games Belief formation
Zhang et al. (zhang2023exploring, ) Collaboration tasks Reflection & debate
GovSim (piatti2024cooperate, ) Resource sustainability Collective cognition
AgentGroupChat (gu2024agent, ) Group deliberation Belief & memory
5.4.1. Building Social Simulacra Mirroring Real-world Society.
With the rapid development of LLM agents, building realistic social simulation systems has become increasingly feasible. A well-known example is AI Town (park2023generative, ), a sandbox environment composed of generative agents that exhibit believable individual behaviors and, at the group level, produce emergent dynamics resembling those in real communities. These systems illustrate how social simulacra can serve as explicit world models, providing environments where societal intelligence—collective sensing, reasoning, and coordination—can be observed and studied.
In social networks, S3 (gao2023s, ) shows that LLM agents can reproduce realistic patterns of information diffusion, capturing the dynamics of public events as they unfold. Papachristou et al. (papachristou2024network, ) further highlight the spontaneous formation of network structures among agents, mirroring the self-organization of human societies. Such works reveal how simulacra embody the adaptive and communicative aspects of societal intelligence in digital form. Beyond networks, LLM agents have demonstrated the capacity to model higher-order reasoning in strategic interactions. Xu et al. (xu2023exploring, ) show that agents in social deduction games like Werewolf display strategic behaviors such as deception and confrontation, reflecting cognitive and competitive dimensions of societal intelligence. In economics and resource allocation, LLM-based agents enable bottom-up modeling of macro-level outcomes from micro-level reasoning. EconAgent (li2024econagent, ) reproduces macroeconomic trends emerging from individual decision-making, SRAP-Agent (ji2024srap, ) evaluates policy effects in resource allocation, and Project Sid (al2024project, ) explores collective responses to taxation rules. These examples illustrate how societal intelligence can manifest as aggregate patterns in economic systems.
More recently, research has shifted toward large-scale platforms that generalize across multiple domains. Among them, AgentSociety (piao2025agentsociety, ) stands out as the most advanced effort to date, providing a large-scale and versatile environment for examining polarization, policy interventions, and other phenomena central to societal intelligence. Building on this direction, platforms such as GenSim (tanggensim, ), YuLan-OneSim (wang2025yulan, ), OASIS (yangoasis, ), and SocioVerse (zhang2025socioverse, ) further extend the vision by integrating diverse social contexts and scaling simulations to thousands or even millions of agents.
Together, these studies demonstrate that LLM-driven social simulacra can serve as explicit world models of human societies, enabling systematic investigations of societal intelligence across social, strategic, and economic domains.
@@ 5.4.2. Agent’s Understanding of External World in Social Simulacra
Beyond mirroring societies at the macro level, social simulacra also enable the study of how agents form internal representations of their environment. Through interaction, LLM agents accumulate experiences, store them as memory, and transform them into implicit world models. These models provide the cognitive substrate for societal intelligence, allowing agents not only to recall past interactions but also to reason about other agents and the broader environment when making decisions (zhang2024survey, ).
Several works demonstrate how implicit world models emerge in practice. Agent-Pro (zhang2024agent, ) transforms interaction histories into structured beliefs, which then guide subsequent decision-making and strategy updates. These beliefs reflect an agent’s understanding of others and connect directly to Theory of Mind capabilities discussed in Section 3.2. Zhang et al. (zhang2023exploring, ) further extend this direction by introducing reflection and debate mechanisms from social psychology to improve collaboration in multi-agent tasks.
At the collective level, GovSim (piatti2024cooperate, ) investigates whether sustainable cooperation can arise in a society of LLM agents. In this setup, each agent gathers information about shared resources and the behavioral strategies of peers through dialogue, forming higher-level insights about the external environment. These insights amount to implicit representations of the world model that underpin group-level societal intelligence. Another application is Interactive Group Chat (gu2024agent, ), which explores human-like deliberation across scenarios such as inheritance disputes and court debates. Here, agents leverage memory and reasoning to produce strategies and social dynamics that closely resemble real human interactions.
5.5. Functions of World Models
At their core, world models are designed to receive external commands or actions and model the dynamic state transitions of an environment. Their functionalities can be broadly categorized into two main roles: acting as cloud-based environments and serving as edge-side agent brains. Cloud-based world models typically manifest as video generation systems, synthesizing large volumes of high-quality video data driven by text or action trajectories. This generated data can function as a data engine, augmenting real-world data for training policy models (e.g., VLA and VLN models). Furthermore, cloud-based world models can act as environments within reinforcement learning, interacting with agents to facilitate evolutionary learning in a virtual setting. This capability significantly reduces the cost and risks associated with real-world interactions, proving especially critical in domains like autonomous driving. Additionally, cloud-based world models can serve as policy evaluators, outputting observation sequences through interaction with policy models, thereby enabling the assessment of policy model performance. In contrast, world models acting as edge-side agent brains often do not require low-level visual generation. Instead, they can compress world states within a latent space; for example, V-JEPA 2 (assran2025v, ) trains a latent space world model that enables on-device action planning through model predictive control. Alternatively, a two-stage approach (hu2024video, ) can be employed, where the world model first processes visual observations, which are then converted into executable actions.
@@ 6.1. Physical Rules and Counterfactual Simulation
A key objective of world models is to capture the causal structure of their environments – especially the underlying physical rules – so they can reason about counterfactuals beyond the data distribution (pearl2009causal, ). This capacity is crucial for handling rare, mission-critical events (e.g., autonomous-driving corner cases (feng2023dense, )) and for narrowing sim-to-real gaps. Recent progress raises the question of whether large-scale, purely data-driven generative models can acquire such rules from raw visual data alone. While transformer- and diffusion-based video generators such as Sora (sora2024, ) produce strikingly realistic sequences, studies reveal persistent physical-law failures – e.g., inaccurate gravity, fluid, or thermal dynamics (wang2023newton, ).
Hybrid approaches that explicitly embed physics are emerging as promising alternatives. Genesis(Genesis, ) illustrates this direction by unifying fast, photo-realistic rendering with a re-engineered universal physics core, allowing language-conditioned data generation grounded in first-principles simulation. PhysGen(liu2024physgen, ) takes a similar stance at the image-to-video level: it couples a rigid-body simulator with a diffusion refiner, enabling controllable, physically plausible motion from a single image. Complementarily, soft-constraint hybrids enforce physics via learning-time priors. physics-informed diffusion introduces PDE-based residual losses that penalize violations of governing equations while preserving generative flexibility (bastek2025physicsinformed, ). Such “hard+soft” designs improve controllability and interpretability without sacrificing realism.
Complementary diagnostic work underscores why such hybrids are needed. Kang et al.(kang2024far, ) show that scaling diffusion video models yields perfect in-distribution fidelity yet breaks down on out-of-distribution or combinatorial tests, indicating “case-based” rather than rule-based generalization. Motamed et al.(motamed2025generative, ) reach a similar conclusion with the Physics-IQ benchmark: current video generators achieve visual realism but largely fail on tasks requiring understanding of optics, fluid dynamics, or magnetism. In parallel, first-principles benchmarks have begun to operationalize “physical fidelity” as a measurable axis: T2VPhysBench (guo2025t2vphysbench, ) evaluates adherence to core physical laws (including Newtonian mechanics and conservation principles) and documents systematic violations across leading text-to-video systems, while VBench-2.0 (zheng2025vbench, ) explicitly introduces Physics and Commonsense as standard evaluation dimensions for video generation.
Taken together, the evidence suggests that data-driven scaling alone is insufficient to recover robust physical laws. Integrating explicit simulators – or otherwise enforcing physical priors (shi2022learning, ) – remains a promising path toward world models that generalize to unseen counterfactual scenarios while retaining interpretability and transparency.